OpenStudy (anonymous):

find the general term an for the geometric sequence a1=−2 and a2=6

OpenStudy (anonymous):

the n is supposed to be lowered as well as the 2 and 1

hartnn (hartnn):

first find common ratio-> r=next term / current term = 6/(-2)=..? then just put it in general formula for n terms \(a_n = a_1 r^{n-1}\) a1 = 1st term =-2